OpenStudy (anonymous):

g(t)=t^3-3+5 what is (g^-1)'(t)

OpenStudy (freckles):

is your g(t) written correctly?

OpenStudy (freckles):

just asknig because as is it is t^3+2

OpenStudy (anonymous):

my bad i meant t^3-3t+5

OpenStudy (misty1212):

HI you are not going to find the inverse

OpenStudy (freckles):

what is the t value here

OpenStudy (freckles):

\[(g^{-1})'(t)=\frac{1}{g'(g^{-1}(t))}\] at what number t

OpenStudy (anonymous):

so i take the inverse and plug in t?

OpenStudy (freckles):

well you aren't going to be able to find the inverse relation as my dear friend @misty1212 said

OpenStudy (anonymous):

why not

OpenStudy (misty1212):

i bet it is \[(g^{-1})'(2)=\frac{1}{g'(g^{-1}(2))}\] or sommat

OpenStudy (misty1212):

because you do not know how to solve a cubic equation

OpenStudy (freckles):

can you solve this for t: y=t^3-3t+5 because I can't

OpenStudy (anonymous):

whered u get 2

OpenStudy (anonymous):

can i just set t^3-3t+5=5

OpenStudy (misty1212):

i made it up

OpenStudy (freckles):

so t is 5

OpenStudy (anonymous):

and that t value is g^-1(5)

OpenStudy (freckles):

\[t^3-3t+5=5\] observe which t satisfies this

OpenStudy (misty1212):

use \(t=0\)

OpenStudy (anonymous):

t= 0 and t = sqrt(3)

OpenStudy (anonymous):

which one should i use

OpenStudy (freckles):

i don't think the latter works

OpenStudy (anonymous):

but t^3-3t=0

OpenStudy (anonymous):

and take out t and its t^2-3=0

OpenStudy (anonymous):

and t=0

OpenStudy (freckles):

oops I was wrong lol this function doesn't seem to be one to one

OpenStudy (freckles):

do you have any restrictions on g

OpenStudy (anonymous):

the question before says find the largest interval on which g has an inverse

OpenStudy (freckles):

that actually gives you 3 values i bet you are going to use the t=0

OpenStudy (anonymous):

containing 0

OpenStudy (freckles):

then 0!

OpenStudy (freckles):

lol

OpenStudy (misty1212):

makes no difference, the derivative still exists locally

OpenStudy (anonymous):

so -1/3

OpenStudy (freckles):

\[(g^{-1})'(0)=\frac{1}{g'(0)}\]

OpenStudy (freckles):

sounds great